### 14 Apr 2026 v1

# PianoFlow: Music-Aware Streaming Piano Motion Generation with Bimanual Coordination

Abstract.
Audio-driven bimanual piano motion generation requires precise modeling of complex musical structures and dynamic cross-hand coordination. However, existing methods often rely on acoustic-only representations lacking symbolic priors, employ inflexible interaction mechanisms, and are limited to computationally expensive short-sequence generation. To address these limitations, we propose PianoFlow, a flow-matching framework for precise and coordinated bimanual piano motion synthesis. Our approach strategically leverages MIDI as a privileged modality during training, distilling these structured musical priors to achieve deep semantic understanding while maintaining audio-only inference. Furthermore, we introduce an asymmetric role-gated interaction module to explicitly capture dynamic cross-hand coordination through role-aware attention and temporal gating. To enable real-time streaming generation for arbitrarily long sequences, we design an autoregressive flow continuation scheme that ensures seamless cross-chunk temporal coherence. Extensive experiments on the PianoMotion10M dataset demonstrate that PianoFlow achieves superior quantitative and qualitative performance, while accelerating inference by over 9× compared to previous methods.
Motion Generation, Multimodal Learning
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYccs: Computing methodologies Motion processing
1. Introduction
Figure 1. PianoFlow enables high-fidelity and real-time bimanual piano motion generation. (Left) Our audio-driven framework distills privileged musical priors and employs Asymmetric Role-Gated Interaction for coordinated bimanual synthesis. (Right) PianoFlow achieves state-of-the-art performance across all metrics, supporting seamless real-time generation.
Audio-driven human motion generation has progressed significantly, enabling applications in VR/AR, digital humans, and robotics (Zhu et al., 2023; Sui et al., 2026; Jiang et al., 2024). This field encompasses well-studied subproblems, such as audio-driven dance (Siyao et al., 2022; Li et al., 2024; Tseng et al., 2023), talking head (Shen et al., 2023; Peng et al., 2023; Richard et al., 2021), and co-speech gesture generation (Zhi et al., 2023). Recently, research has expanded to instrument performance synthesis, exploring tasks such as violin (Nishizawa et al., 2025; Jin et al., 2024), Guzheng (Chen et al., 2021), and piano (Gan et al., 2024) motion generation. Unlike conventional tasks that rely on coarse rhythmic or phonetic alignment, instrument performance demands precise, high-frequency finger articulation, strict physical interaction constraints, and deep alignment with structured musical harmony. In this work, we focus on the challenging task of audio-driven bimanual piano motion generation, tackling the unique challenge of dynamic cross-hand coordination across a dense spatial keyboard.
Early works like PianoPlayer (Musy, 2018) generate static hand poses from piano scores, relying on classifiers to predict fingering. Subsequently, BACH (Jiao et al., 2025) transforms scores into continuous hand motion using a two-stage deep generative framework that predicts keyframe guidance and synthesizes smooth transitions. Driven by the emergence of aligned datasets, recent research has shifted toward audio-conditioned generation. PianoMotion (Gan et al., 2024) leverages diffusion models to generate continuous sequences from audio, while S2C (Liu et al., 2025c) decouples hands and models them separately, followed by a static interaction mechanism.
However, synthesizing realistic bimanual motion directly from continuous audio remains challenging due to three primary limitations in existing methods: (1) Limitations of acoustic-only representations: Existing methods often rely on speech-based encoders that lack music-specific sensitivity. Furthermore, without explicit symbolic priors, they struggle to capture fine-grained note onsets and harmonic structures from complex audio. (2) Inflexible bimanual coordination: The hands assume distinct, time-varying roles (e.g., melody vs. accompaniment). Static rules or weak coupling fail to capture these dynamic shifts, often resulting in bimanual interference. (3) Limited scalability and real-time efficiency: Most frameworks are computationally intensive and restricted to fixed-length sequences, precluding streaming deployment for arbitrarily long performances.
To address these challenges, we propose PianoFlow, a two-stage flow-matching framework as illustrated in Figure 1. The system comprises Music-Aware Wrist Trajectory Generation and Bimanual Coordinated Gesture Generation. Unlike prior works using speech encoders like Wav2Vec2.0 (Baevski et al., 2020) or Hubert (Hsu et al., 2021), we adopt MuQ (Zhu et al., 2025) for music-specific representations. Furthermore, we leverage MIDI (Rothstein, 1995) as a privileged modality, employing a Harmonic Perceiver to capture octave and interval relationships. By distilling a multimodal teacher into an audio-only student, we inject MIDI-level structural priors into the learned representations while maintaining audio-driven inference (Mansourian et al., 2025; Aslam et al., 2023). Additionally, we adopt a decoupled bimanual formulation while enabling interaction at the feature bottleneck. We propose an Asymmetric Role-Gated Interaction module (ARGI), which utilizes role-specific biases as identity cues to facilitate information exchange. A temporal gating mechanism further modulates interaction strength, enabling adaptive coordination while reducing interference during role divergence.
[…]
Extensive experiments on PianoMotion10M, encompassing both short- and long-sequence generation, demonstrate the superiority of our architecture through quantitative and qualitative comparisons. Furthermore, ablation studies investigate the specific role of each individual component. These results confirm that PianoFlow effectively interprets musical semantics to synthesize precise and coordinated playing motions.
The main contributions of this work are summarized as follows:
  • We propose PianoFlow, a two-stage framework for audio-driven bimanual motion generation. By leveraging MIDI as a privileged modality, it distills structural priors to achieve a deep understanding of musical semantics while maintaining audio-driven inference.
  • We design an Asymmetric Role-Gated Interaction mechanism to explicitly model bimanual dynamics. By incorporating role-specific biases and temporally modulating interaction strength, this design enables fine-grained and adaptive coordination between the two hands.
  • We develop Autoregressive Flow Continuation, ensuring cross-chunk temporal coherence and real-time, arbitrary-length streaming inference by anchoring to the optimal transport path.
  • Extensive experiments on the PianoMotion10M dataset demonstrate that PianoFlow achieves superior performance in both quantitative and qualitative evaluations, while accelerating inference by over $9\times$ compared to prior methods.
2. Related Work
2.1. Audio-Driven Motion Generation
Recent advances in generative modeling have significantly improved audio-driven human motion synthesis across diverse applications, including co-speech motion generation (Shen et al., 2023; Sun et al., 2024; Aneja et al., 2024; Cassell et al., 2001; Zhi et al., 2023; Liu et al., 2024b; Yi et al., 2023; Chu et al., 2025; Jafari et al., 2024; Liu et al., 2024a; Xu et al., 2024; Chhatre et al., 2024; Chen et al., 2025b; Fan et al., 2024; Kim et al., 2025a; Xu et al., 2026; Liu et al., 2022b; Chae-Yeon et al., 2025; Kim et al., 2025b; Liu et al., 2025b), music-driven dance generation (Siyao et al., 2022; Tseng et al., 2023; Li et al., 2023; Yang et al., 2025b; Li et al., 2024; Guo et al., 2025; Dai et al., 2026; Liu et al., [n. d.], 2025a; Ghosh et al., 2025; Yang et al., 2025c; Li et al., 2025; Wang et al., 2025a; Fan et al., 2025; Huang et al., 2024), and specialized scenarios (Xu et al., 2025). Early methods focused on learning direct mappings from audio to motion using adversarial or variational frameworks (Kim et al., 2022). In contrast, recent approaches increasingly leverage large-scale generative architectures, such as Transformers and diffusion models, to capture complex spatio-temporal dynamics (Sun et al., 2024). To enhance motion realism and temporal coherence, many strategies have been proposed. For instance, several pioneering works employ explicit structural representations to achieve accurate synchronization and expressive generation (Richard et al., 2021; Zhang et al., 2023). Transformer-based models have further enhanced the modeling of long-range temporal dependencies (Fan et al., 2022), while diffusion-based approaches have demonstrated the ability to synthesize expressive motion sequences directly from audio without requiring heavy intermediate structural constraints (Wei et al., 2024; Liu et al., 2025b). However, synthesizing realistic piano motions requires parsing complex harmonic structures and fine-grained note onsets, a rigorous challenge absent in general body dynamics. The reliance of existing models on coarse-grained acoustic features limits their ability to meet the stringent demands of precise finger articulation. To bridge this semantic gap, we introduce a cross-modal distillation scheme to embed explicit musical priors into the audio representation, establishing a highly precise semantic foundation tailored for piano performance.
2.2. Piano Hand Motion Generation
Piano hand motion generation requires the precise modeling of fine-grained finger articulation and coordinated bimanual interaction. Existing approaches generally follow two paradigms: embodied robotic control and data-driven kinematic synthesis. The former focuses on learning dexterous manipulation within physical simulators via reinforcement or imitation learning (Zakka et al., 2023; Zeulner et al., 2025). For instance, RoboPianist (Zakka et al., 2023) established a physics-based framework leveraging deep RL and fingering priors for complex bimanual performances, while FürElise (Wang et al., 2024) integrated generative modeling by guiding a physics-based controller with diffusion-generated kinematic trajectories. The latter paradigm synthesizes hand kinematics directly from musical signals, supporting expressive applications such as AI-assisted education and digital human animation. Within this data-driven setting, Gan et al. (Gan et al., 2024) introduced the PianoMotion10M dataset alongside a diffusion-based baseline, while Liu et al. (Liu et al., 2025c) modeled bimanual coordination through a dual-stream diffusion architecture. However, these methods rely on static interactions that fail to capture dynamic, asymmetric cross-hand role shifts (e.g., melody vs. accompaniment). Furthermore, their diffusion pipelines suffer from prohibitive sampling latency and are restricted to short-sequence generation. To overcome these bottlenecks, PianoFlow introduces an Asymmetric Role-Gated Interaction (ARGI) module to explicitly model time-varying bimanual dependencies, alongside an Autoregressive Flow Continuation (AFC) mechanism for seamless, real-time streaming synthesis of arbitrary-length sequences.
Figure 2. Overview of the PianoFlow architecture. The framework operates in two stages: (1) Music-Aware Wrist Trajectory Generation utilizes a decoupled cross-modal distillation scheme to transfer structured harmonic knowledge from a multimodal teacher (MIDI and audio) to an audio-only student. (2) Bimanual Coordinated Gesture Generation synthesizes fine-grained kinematics via conditional flow-matching. Specifically, a factorized spatio-temporal backbone estimates the velocity field, while an Asymmetric Role-Gated Interaction (ARGI) module explicitly models bimanual coordination at the network bottleneck.
3. Method
Our goal is to synthesize precise and well-coordinated bimanual piano motions directly from audio. As illustrated in Fig. 2, the proposed PianoFlow framework consists of two stages: (1) Music-Aware Wrist Trajectory Generation and (2) Bimanual Coordinated Gesture Generation. In the first stage, we introduce a cross-modal distillation scheme that leverages MIDI as a privileged modality, transferring structured musical priors to the audio-driven student model. In the second stage, we employ a spatio-temporal flow-matching framework to model fine-grained finger kinematics, incorporating an Asymmetric Role-Gated Interaction (ARGI) mechanism to explicitly capture dynamic cross-hand coordination. For real-time inference, we adopt an Autoregressive Flow Continuation (AFC) strategy, enabling coherent streaming synthesis for sequences of arbitrary length.
3.1. Music-Aware Wrist Trajectory Generation
Given an input audio sequence, this stage predicts bimanual wrist trajectories $\varphi=[\varphi^{L},\varphi^{R}]\in\mathbb{R}^{N\times 6}$, where $N$ denotes the sequence length. To explicitly ground these trajectories in musical semantics, a cross-modal distillation scheme transfers structured harmonic priors from a multimodal (MIDI and audio) teacher to an audio-only student. Considering the distinct behaviors of each hand, we model their trajectories using independent networks with identical architectures but unshared parameters. This decoupled design prevents premature coupling, allowing each branch to capture hand-specific kinematic priors.
3.1.1. Decoupled Cross-Modal Architecture.
We adopt a decoupled teacher-student architecture, where modality-specific features are learned independently and aligned in a shared latent space.
Student Branch Encoding. Taking raw audio as input, we extract acoustic features using a frozen MuQ (Zhu et al., 2025) encoder, which is pre-trained on large-scale music datasets for robust music representation. These features are then processed through a Student Encoder to yield $\mathbf{F}_{\sigma}^{\text{enc}}$.
Teacher Branch Encoding. The teacher branch leverages MIDI as a privileged modality to provide explicit symbolic priors. We design a Harmonic Perceiver to encode pitch-wise dependencies from the piano-roll input $\mathcal{X}_{\text{midi}}\in\mathbb{R}^{N\times 88}$, where each frame represents note activations across the 88 standard piano keys.
To capture the hierarchical nature of musical intervals within each frame $n$, we model the local harmonic structure along the pitch dimension using a cascaded convolutional design: an octave-aware kernel (size $13$) followed by an interval-aware kernel (size $5$). To retain the spatial keyboard topology, we explicitly inject an absolute pitch positional encoding $\mathbf{P}_{\text{pitch}}\in\mathbb{R}^{88\times C}$. This yields a pitch-wise structural representation:
(1) $\mathbf{S}_{n}=\psi_{\text{int}}\left(\psi_{\text{oct}}(\mathcal{X}_{\text{midi},n})\right)+\mathbf{P}_{\text{pitch}},$
where $\mathbf{S}_{n}\in\mathbb{R}^{88\times C}$.
To aggregate this harmonic information, inspired by the asymmetric attention mechanism of the Perceiver (Jaegle et al., 2021), we introduce a set of learnable queries to perform attention-based pooling over the pitch dimension:
(2) $\mathbf{Q}^{p}=\mathbf{Q}_{0}\mathbf{W}_{Q}^{p},\quad\mathbf{K}_{n}^{p}=\mathbf{S}_{n}\mathbf{W}_{K}^{p},\quad\mathbf{V}_{n}^{p}=\mathbf{S}_{n}\mathbf{W}_{V}^{p},$
(3) $\tilde{\mathbf{E}}_{n}=\text{Softmax}\left(\frac{\mathbf{Q}^{p}(\mathbf{K}_{n}^{p})^{\top}}{\sqrt{d_{p}}}\right)\mathbf{V}_{n}^{p},$
where $\mathbf{Q}_{0}\in\mathbb{R}^{M\times C}$ represents a set of $M$ learnable queries, $\mathbf{W}_{Q}^{p},\mathbf{W}_{K}^{p},\\ \mathbf{W}_{V}^{p}$ are learnable projection matrices, and $d_{p}$ denotes the hidden dimension of the pitch keys. This attention mechanism enables the model to selectively focus on distinct harmonic regions (e.g., melody and bass). Subsequently, we map $\tilde{\mathbf{E}}_{n}$ into a unified embedding $\mathbf{E}_{n}$. Finally, a temporal self-attention encoder processes the sequence $\{\mathbf{E}_{n}\}_{n=1}^{N}$ to yield the final MIDI representation $\mathbf{F}_{\text{midi}}$.
To fuse these modalities, the Teacher Encoder conditions the acoustic representation $\mathbf{F}_{\sigma}^{\text{enc}}$ on the aggregated MIDI features via cross-attention. Specifically, the acoustic features act as queries, while the MIDI features serve as keys and values:
(4) $\mathbf{Q}^{a}=\mathbf{F}_{\sigma}^{\text{enc}}\mathbf{W}_{Q}^{a},\quad\mathbf{K}^{m}=\mathbf{F}_{\text{midi}}\mathbf{W}_{K}^{m},\quad\mathbf{V}^{m}=\mathbf{F}_{\text{midi}}\mathbf{W}_{V}^{m},$
(5) $\tilde{\mathbf{F}}=\text{Softmax}\left(\frac{\mathbf{Q}^{a}(\mathbf{K}^{m})^{\top}}{\sqrt{d_{m}}}\right)\mathbf{V}^{m},$
(6) $\mathbf{F}_{\tau}^{\text{enc}}=\mathbf{F}_{\sigma}^{\text{enc}}+\tilde{\mathbf{F}}$,
where $\mathbf{W}_{Q}^{a},\mathbf{W}_{K}^{m},\mathbf{W}_{V}^{m}$ are learnable projection matrices mapping the features to a joint attention space of dimension $d_{m}$. This operation explicitly injects structured musical priors into the acoustic representation, yielding the final teacher encoding $\mathbf{F}_{\tau}^{\text{enc}}$.
Shared Backbone and Decoupled Decoding. We process the encoded student and teacher features, $\mathbf{F}_{\sigma}^{\text{enc}}$ and $\mathbf{F}_{\tau}^{\text{enc}}$, through a Transformer backbone shared across the student and teacher branches to map them into a unified latent space. Two decoupled decoding heads then generate $\mathbf{F}_{\sigma}^{\text{dec}}$ and $\mathbf{F}_{\tau}^{\text{dec}}$, respectively. A linear projection further maps each decoded feature to wrist trajectories, yielding $\varphi_{\sigma},\varphi_{\tau}\in\mathbb{R}^{N\times 6}$. This design explicitly aligns the high-level representations within the shared space while preventing the student from directly fitting the teacher’s exact output distribution.
3.1.2. Cross-Modal Distillation.
We achieve cross-modal distillation by aligning features at both the encoder and decoder stages. To ensure temporal consistency, we formulate the alignment loss by minimizing the cosine distance across the sequence dimension:
(7) $\mathcal{L}_{\text{align}}=\sum_{l\in\{\text{enc},\text{dec}\}}\frac{1}{N}\sum_{n=1}^{N}\left(1-\frac{\langle\mathbf{F}_{\sigma,n}^{l},\text{sg}(\mathbf{F}_{\tau,n}^{l})\rangle}{\|\mathbf{F}_{\sigma,n}^{l}\|_{2}\,\|\text{sg}(\mathbf{F}_{\tau,n}^{l})\|_{2}}\right),$
where $\text{sg}(\cdot)$ denotes the stop-gradient operator.
While the MuQ acoustic encoder is pre-trained, the Harmonic Perceiver is trained from scratch. Consequently, enforcing strong alignment during early training may introduce unreliable supervision from the under-trained teacher. To mitigate this, we progressively increase the distillation loss weight $\lambda_{\text{distill}}$ during training. This strategy enables the model to learn robust modality-specific representations before enforcing strict cross-modal consistency.
Each branch is supervised by a task loss that enforces accurate position reconstruction and temporal smoothness. Let $\varphi\in\mathbb{R}^{N\times 6}$ denote the ground-truth trajectory, and $\hat{\varphi}_{b}$ denote the predicted trajectory from branch $b\in\{\sigma,\tau\}$. The task loss is defined as:
(8) $\mathcal{L}_{\text{task}}^{b}=\mathcal{L}_{\text{rec}}^{b}+\lambda_{\text{vel}}\mathcal{L}_{\text{vel}}^{b},$
where the reconstruction loss $\mathcal{L}_{\text{rec}}^{b}$ and velocity loss $\mathcal{L}_{\text{vel}}^{b}$ are formulated as:
(9) $\mathcal{L}_{\text{rec}}^{b}=\frac{1}{N}\sum_{n=1}^{N}\|\hat{\varphi}_{b,n}-\varphi_{n}\|_{2}^{2},$
(10) $\mathcal{L}_{\text{vel}}^{b}=\frac{1}{N-1}\sum_{n=1}^{N-1}\|(\hat{\varphi}_{b,n+1}-\hat{\varphi}_{b,n})-(\varphi_{n+1}-\varphi_{n})\|_{2}^{2}.$
Here, $\|\cdot\|_{2}^{2}$ denotes the squared $L_{2}$ norm, and $\lambda_{\text{vel}}$ is a hyperparameter controlling temporal smoothness.
The overall objective is:
(11) $\mathcal{L}_{\text{total}}=\mathcal{L}_{\text{task}}^{\sigma}+\lambda_{\tau}\mathcal{L}_{\text{task}}^{\tau}+\lambda_{\text{distill}}\mathcal{L}_{\text{align}},$
where $\lambda_{\tau}$ and $\lambda_{\text{distill}}$ control the contributions of the teacher supervision and distillation alignment, respectively. Ultimately, this stage effectively distills structured musical priors into the audio-driven student model while maintaining an efficient audio-only inference pipeline.
3.2. Bimanual Coordinated Gesture Generation
This stage synthesizes fine-grained bimanual hand gestures $\mathbf{x}=[\mathbf{x}^{L},\mathbf{x}^{R}]\in\mathbb{R}^{N\times D}$. As illustrated in Figure 2, we adopt a flow-matching framework (Lipman et al., 2022), where a dual-stream spatio-temporal architecture independently predicts the continuous-time velocity fields for each hand. To bridge these unshared streams, a parameter-shared Asymmetric Role-Gated Interaction (ARGI) module explicitly models dynamic cross-hand coordination at the network bottleneck. During inference, we sample gestures by solving the Ordinary Differential Equation (ODE) induced by the learned velocity field.
3.2.1. Velocity Field Network.
The Velocity Field Network serves as the core generative backbone, designed to map the intermediate gesture state and conditioning features to a continuous-time velocity field.
Conditioning. We construct a global conditioning feature $\mathbf{c}=[\mathbf{F}_{\text{audio}},\mathbf{F}_{\sigma}^{\text{enc}},\hat{\varphi}_{\sigma}]$, where $\mathbf{F}_{\text{audio}}$ denotes the frozen MuQ acoustic representation, $\mathbf{F}_{\sigma}^{\text{enc}}$ represents the student-side encoded feature, and $\hat{\varphi}_{\sigma}$ is the predicted bimanual wrist trajectory from Stage 1. At time $t\in[0,1]$, the intermediate latent gesture state is denoted by $\mathbf{x}_{t}=[\mathbf{x}_{t}^{L},\mathbf{x}_{t}^{R}]$.
Factorized Spatio-Temporal Backbone. Specifically, a Spatial Transformer first models the intra-hand structural correlations by tokenizing individual kinematic joints to perform exhaustive intra-hand self-attention. Subsequently, a temporal 1D U-Net (Ronneberger et al., 2015) processes these aggregated spatial embeddings to capture multi-scale gesture dynamics along the temporal axis. The global conditioning $\mathbf{c}$ is injected into the intermediate features of each U-Net block via Feature-wise Linear Modulation (Perez et al., 2018). At the network bottleneck, we extract the compressed representations $\mathbf{h}_{s}\in\mathbb{R}^{N^{\prime}\times C}$ for each hand $s\in\{L,R\}$, serving as the foundation for the subsequent cross-hand interaction. Ultimately, the network outputs the conditional velocity field $\mathbf{v}_{\theta}(t,\mathbf{x}_{t};\mathbf{c})$ for the concatenated bimanual representation.
Asymmetric Role-Gated Interaction (ARGI). To explicitly model coordinated bimanual dynamics, we integrate cross-hand dependencies at the U-Net bottleneck. Given the bottleneck features $\mathbf{h}_{s}$, let $\bar{s}$ represent the contralateral hand. Unlike the branch-specific backbones, ARGI serves as a parameter-shared bridge to facilitate information exchange. To instill hand-specific awareness, the query $\mathbf{Q}_{s}^{\mathcal{I}}$ is derived by injecting a learnable hand identity embedding $\mathbf{e}_{\text{id}}(s)$ followed by a self-attention layer $\zeta_{\text{self}}$. Crucially, we introduce learnable projection matrices $\mathbf{W}_{Q}^{\mathcal{I}},\mathbf{W}_{K}^{\mathcal{I}},\mathbf{W}_{V}^{\mathcal{I}}\in\mathbb{R}^{C\times d_{k}}$ to map the representations into a dedicated interaction space:
(12) $\mathbf{Q}_{s}^{\mathcal{I}}=\zeta_{\text{self}}(\mathbf{h}_{s}+\mathbf{e}_{\text{id}}(s))\mathbf{W}_{Q}^{\mathcal{I}},\quad\mathbf{K}_{\bar{s}}^{\mathcal{I}}=\mathbf{h}_{\bar{s}}\mathbf{W}_{K}^{\mathcal{I}},\quad\mathbf{V}_{\bar{s}}^{\mathcal{I}}=\mathbf{h}_{\bar{s}}\mathbf{W}_{V}^{\mathcal{I}}.$
The cross-hand message $\mathbf{m}_{s}$ is then synthesized via a cross-attention mechanism:
(13) $\mathbf{m}_{s}=\mathrm{Softmax}\left(\frac{\mathbf{Q}_{s}^{\mathcal{I}}(\mathbf{K}_{\bar{s}}^{\mathcal{I}})^{\top}}{\sqrt{d_{k}}}\right)\mathbf{V}_{\bar{s}}^{\mathcal{I}},$
where $d_{k}$ denotes the channel dimension of the key space. To accommodate the dynamic nature of piano performance (e.g., shifts between melody and accompaniment roles), we employ a temporal gating mechanism to modulate the interaction strength. Conditioned on the local context of $\mathbf{h}_{s}$, a multi-layer perceptron (MLP) predicts a frame-wise gate $\mathbf{g}_{s}\in[0,1]^{N^{\prime}}$. The cross-hand message is then gated, refined by a feed-forward network (FFN), and integrated into the main flow:
(14) $\mathbf{g}_{s}=\sigma\big(\mathrm{MLP}(\mathbf{h}_{s})\big),\quad\mathbf{h}^{\text{inter}}_{s}=\mathbf{h}_{s}+\mathrm{FFN}\big(\mathbf{g}_{s}\odot\mathbf{m}_{s}\big),$
where $\sigma(\cdot)$ is the sigmoid function and $\odot$ denotes element-wise multiplication. The updated features $\mathbf{h}^{\text{inter}}_{s}$ are subsequently processed by the U-Net decoder to estimate the final velocity field. This design enables adaptive bimanual coordination by dynamically evaluating the implicit musical role of each hand at each time step.
3.2.2. Flow Matching and Training Objective.
To optimize the Velocity Field Network, we employ a conditional flow matching (CFM) framework, which provides a simulation-free regression objective for learning the continuous-time generative process.
Conditional Flow Formulation. We learn a continuous-time conditional velocity field $\mathbf{v}_{\theta}(t,\mathbf{x}_{t};\mathbf{c})$ that transports samples from a base distribution to the target data distribution, conditioned on the multi-modal features $\mathbf{c}$. During training, we draw a ground-truth gesture sample $\mathbf{x}_{1}\sim p_{\text{data}}$, a noise sample $\mathbf{x}_{0}\sim p_{\text{base}}$ from a standard Gaussian $\mathcal{N}(\mathbf{0},\mathbf{I})$, and a time step $t\sim\mathcal{U}(0,1)$.
Following Optimal Transport (OT) principles (Lipman et al., 2022; Liu et al., 2022a), we define a linear probability path:
(15) $\mathbf{x}_{t}=(1-t)\mathbf{x}_{0}+t\mathbf{x}_{1},$
which induces a constant target velocity $\mathbf{u}_{t}=\mathbf{x}_{1}-\mathbf{x}_{0}$. The network is trained to estimate the velocity field $\mathbf{v}_{\theta}(t,\mathbf{x}_{t};\mathbf{c})$ by minimizing the conditional flow-matching objective:
(16) $\mathcal{L}_{\mathrm{FM}}=\mathbb{E}_{\mathbf{x}_{0},\mathbf{x}_{1},t}\left[\left\|\mathbf{v}_{\theta}(t,\mathbf{x}_{t};\mathbf{c})-\mathbf{u}_{t}\right\|_{2}^{2}\right].$
During training, the ARGI module facilitates explicit cross-hand coupling within the network’s intermediate layers, ensuring the synthesis of tightly coordinated gesture flows.
At inference time, gestures are generated by sampling an initial noise $\mathbf{x}_{0}\sim p_{\text{base}}$ and solving the following Ordinary Differential Equation (ODE) from $t=0$ to $t=1$ using numerical solvers (Chen et al., 2018):
(17) $\frac{\mathrm{d}\mathbf{x}(t)}{\mathrm{d}t}=\mathbf{v}_{\theta}\big(t,\mathbf{x}(t);\mathbf{c}\big),\qquad\mathbf{x}(0)=\mathbf{x}_{0}.$
While standard adaptive-step ODE solvers are effective for fixed-length sequences, this formulation lacks the temporal continuity required for streaming synthesis of arbitrary length.
3.3. Streaming Inference via Autoregressive Flow Continuation
For arbitrary-length motion generation, we propose Autoregressive Flow Continuation (AFC). This method uses a causal sliding-window strategy to process gesture chunks sequentially. In this framework, the prior sequence provides temporal grounding for the current chunk via guided continuous-time inpainting in an overlap. To ensure a deterministic transport path, we fix a noise sample $\boldsymbol{\epsilon}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$. We denote the gesture from the preceding chunk’s tail as $\mathbf{x}_{\text{past}}$, anchoring temporal continuity. During forward-time integration from $t$ to $t+\Delta t$, we obtain a provisional update $\mathbf{x}_{t+\Delta t}^{\text{uncond}}$ via a numerical ODE solver step:
(18) $\mathbf{x}_{t+\Delta t}^{\text{uncond}}=\Psi(\mathbf{x}_{t},\mathbf{v}_{\theta},t,\Delta t,\mathbf{c}),$
where $\Psi$ denotes the solver’s transition function. Simultaneously, we construct a reference state $\mathbf{x}_{t+\Delta t}^{\text{known}}$ for the overlapping region by interpolating the fixed noise $\boldsymbol{\epsilon}$ with the known history $\mathbf{x}_{\text{past}}$ along the optimal transport path:
(19) $\mathbf{x}_{t+\Delta t}^{\text{known}}=(1-(t+\Delta t))\boldsymbol{\epsilon}+(t+\Delta t)\mathbf{x}_{\text{past}}.$
We then refine the state by blending the known component with the provisional estimate using a temporal cosine mask $\mathbf{m}$ with values in $[0,1]$:
(20) $\mathbf{x}_{t+\Delta t}\leftarrow\mathbf{m}\odot\mathbf{x}_{t+\Delta t}^{\text{known}}+(\mathbf{1}-\mathbf{m})\odot\mathbf{x}_{t+\Delta t}^{\text{uncond}},$
where $\mathbf{m}$ decays from $1$ to 0 across the overlap region. This mechanism enforces boundary consistency while smoothing the transition into free generation. Executing this update at every solver iteration ensures that each new chunk integrates seamlessly with the prior gesture trajectory, maintaining long-term temporal coherence without inducing additional latency.
Table 1. Quantitative evaluation on the PianoMotion10M dataset under the standard short-sequence setting.
[…]
EmoTalk (Peng et al., 2023) ICCV 2023 0.445 0.232 0.044 0.353 0.360 0.259 0.033 0.313 4.645 -
LivelySpeaker (Zhi et al., 2023) ICCV 2023 0.538 0.220 0.038 0.406 0.535 0.249 0.030 0.334 4.157 -
PianoMotion (Gan et al., 2024) ICLR 2025 0.372 0.217 0.037 0.248 0.351 0.244 0.030 0.205 3.281 1.746
S2C (Liu et al., 2025c) MM 2025 0.309 0.215 0.035 0.229 0.293 0.242 0.028 0.198 3.012 5.723
PianoFlow (Ours) - 0.266 0.210 0.029 0.200 0.263 0.238 0.027 0.184 2.674 0.186
Table 2. Quantitative comparison of piano motion generation under the long-sequence setting on the PianoMotion10M dataset.
Method Left Hand Right Hand
[…]
PianoMotion (Gan et al., 2024) + SWF 0.432 0.220 0.045 0.253 0.334 0.424 0.250 0.037 0.208 0.269
S2C (Liu et al., 2025c) + SWF 0.443 0.223 0.041 0.251 0.350 0.397 0.252 0.035 0.206 0.320
PianoFlow + SWF 0.317 0.223 0.040 0.236 0.324 0.296 0.251 0.032 0.203 0.272
PianoFlow + AFC 0.309 0.222 0.040 0.232 0.312 0.290 0.249 0.032 0.199 0.255
4. Experiments
4.1. Experiments Setting
4.1.1. Datasets.
To validate the effectiveness of PianoFlow, we conduct experiments on the PianoMotion10M dataset (Gan et al., 2024), comprising 1,966 piano performance sequences at 30 FPS, totaling 116 hours. The dataset features synchronized audio, MIDI, and hand motion annotations, comprising 3D wrist coordinates and hand poses parameterized by the MANO (Romero et al., 2017) model. The raw data is partitioned into 30-second clips, yielding 7,519 training, 821 validation, and 8,399 test samples. In line with the configuration of Gan et al. (Gan et al., 2024), we further extract 8-second segments for training and short-sequence evaluation.
4.1.2. Evaluation Metrics.
We evaluate the synthesized motions across three dimensions, including distributional fidelity, kinematic accuracy, and computational efficiency. This evaluation incorporates five base metrics (Gan et al., 2024) along with two extended metrics. (1) Fréchet Inception Distance (FID) quantifies the global distributional discrepancy between synthesized motions and ground-truth data within a pre-trained latent space. (2) Fréchet Gesture Distance (FGD) assesses hand-specific pose distributions by calculating the Fréchet distance within the MANO (Romero et al., 2017) parameter or joint coordinate space. (3) Wasserstein Gesture Distance (WGD) evaluates distributional alignment via optimal transport, calculating the Earth Mover’s Distance (EMD) following PCA and Gaussian Mixture Model (GMM) fitting. (4) Position Distance (PD) quantifies wrist spatial accuracy via the mean Euclidean distance between predicted and ground-truth wrist coordinates across the sequence. (5) Smoothness evaluates physical plausibility by measuring acceleration discrepancies relative to the ground truth to ensure temporal stability. (6) Real-Time Factor (RTF) measures computational efficiency, where an $\text{RTF}<1.0$ indicates real-time generation capability. (7) Frequency Domain Error (FDE) evaluates preservation of high-frequency kinematic details and penalizes over-smoothing artifacts from interpolation-based fusion. We compute amplitude spectra $S(f)$ and $\hat{S}(f)$ for the sequences via Fast Fourier Transform. Defining $\mathcal{H}$ as high-frequency bins exceeding $5$ Hz, FDE is the Mean Absolute Error (MAE) between spectral components to ensure precise motion alignment:
(21) $\text{FDE}=\frac{1}{|\mathcal{H}|}\sum_{f\in\mathcal{H}}\left|S(f)-\hat{S}(f)\right|.$
A lower FDE signifies superior retention of rapid musical articulations and spectral fidelity, ensuring synthesized motions are both energetic and rhythmically accurate.
For long-sequence evaluation, we implement a sliding-window protocol (window size of 240 frames, stride of 120) for FGD, WGD, PD, and Smoothness, reporting the averaged results across all windows.
4.1.3. Baselines.
We benchmark PianoFlow against several state-of-the-art baselines, encompassing both general-purpose motion synthesis frameworks and specialized piano-specific models. (1) EmoTalk (Peng et al., 2023) employs a cross-reconstruction mechanism to learn disentangled latent spaces and utilizes a Transformer-based decoder for sequence generation. (2) LivelySpeaker (Zhi et al., 2023) is a two-stage framework that leverages pre-trained embeddings for conditioning and a diffusion-based refiner to synthesize expressive motions. (3) PianoMotion (Gan et al., 2024) adopts a DDPM-based diffusion framework to jointly model bimanual motions. It generates initial trajectories as coarse guidance to condition a subsequent diffusion-based module for fingering synthesis. (4) S2C (Liu et al., 2025c) builds upon the framework of PianoMotion by adopting a dual-stream diffusion architecture to model bimanual motions independently. It introduces static interaction between the hands during the generation process. Following Gan et al. (Gan et al., 2024), EmoTalk and LivelySpeaker were redeployed and evaluated on the PianoMotion10M dataset.
To facilitate a consistent comparison in long-sequence synthesis, we extend these fixed-length baselines via a Sliding-Window Fusion (SWF) strategy. In this setup, we generate 8-second segments with a 1-second temporal overlap and apply linear cross-fading in the overlapping regions to ensure continuity across chunks.
4.1.4. Implementation Details.
In the first stage, the decoupled cross-modal distillation scheme is trained for 100K iterations with a batch size of 24, using loss weights $\lambda_{\text{vel}}=3.0$, $\lambda_{\tau}=1.0$, and $\lambda_{\text{distill}}=0.1$ to transfer harmonic knowledge from the multimodal teacher to the audio-only student. In the second stage, the conditional flow-matching framework is trained for 250K iterations with a batch size of 16. For both stages, we use 8-second sequence segments and the AdamW optimizer with an initial learning rate of $5\times 10^{-5}$. A ReduceLROnPlateau scheduler (factor 0.5) is applied, with patience values of 10 and 15 for the first and second stages, respectively. During inference, motion is generated using a second-order Heun solver (Karras et al., 2022) with 25 sampling steps. All experiments are conducted on a single NVIDIA RTX 4090 GPU.
Figure 3. Qualitative comparison of synthesized motions, with red and green boxes highlighting kinematic inaccuracies and precise alignments, respectively. Overall, PianoFlow achieves superior spatial accuracy and bimanual coordination.
4.2. Quantitative Results
[…]
4.3. Qualitative Results
We qualitatively compare PianoFlow with competitive baselines, PianoMotion (Gan et al., 2024) and S2C (Liu et al., 2025c). As illustrated in Figure 3, PianoFlow generates more precise wrist trajectories and hand poses for accurate keystroke alignment. Specifically, Figure 3(a) showcases a challenging wide-stretch gesture that our model synthesizes with higher fidelity compared to the baselines. Notably, Figure 3(d) highlights PianoFlow’s ability to capture the inherent flexibility of fingering strategies. Instead of rote memorization, the model explores diverse yet functionally equivalent fingering patterns to strike the same target keys. By synthesizing various biologically plausible poses in strict alignment with the audio, PianoFlow demonstrates its mastery of the keyboard’s underlying physical constraints. Furthermore, our generated sequences maintain high temporal smoothness with minimal jitter throughout continuous performances, as demonstrated in the supplementary material.
4.4. Ablation Study
To evaluate the contribution of individual components, we categorize our ablation models into three dimensions: (1) Music Representations: w/ HuBERT (replacing MuQ with HuBERT) and w/o MIDI (removing MIDI distillation). (2) Gesture Modeling: w/o Decoupling (joint bimanual modeling in Stage 1), w/o Stage 1 (removing wrist trajectory guidance), and w/o Spatial (omitting spatial Transformers). (3) Interaction Mechanisms: w/o ARGI (independent hand modeling) and w/ Cross-Attn (replacing ARGI with vanilla cross-attention). As shown in Table 3, PianoFlow achieves the optimal balance across all metrics. The FID degradation in w/ HuBERT and w/o MIDI confirms that music-specific encoders and the use of a privileged modality are vital for capturing musical semantics. For gesture modeling, w/o Stage 1 causes the left-hand PD to surge from 0.029 to 0.095, demonstrating that wrist trajectories serve as essential kinematic anchors. Although w/o Spatial achieves slightly better distributional alignment in terms of WGD, its Smoothness and FID noticeably degrade. These results and the performance drop in w/o Decoupling collectively justify our architecture design for preserving hand priors and joint correlations. Finally, ARGI outperforms independent modeling (w/o ARGI) and standard cross-attention (w/ Cross-Attn), validating that dynamic temporal gating effectively facilitates bimanual coordination while mitigating interference.
Table 3. Ablation Study on the PianoMotion10M dataset under the standard short-sequence setting.
[…]
PianoFlow 0.266 0.210 0.029 0.200 0.263 0.238 0.027 0.184 2.674
   w/ Hubert 0.293 0.214 0.032 0.251 0.286 0.240 0.029 0.196 2.868
   w/o MIDI 0.277 0.211 0.030 0.203 0.281 0.241 0.028 0.194 2.733
   w/o Decoupling 0.283 0.211 0.038 0.221 0.293 0.241 0.035 0.198 2.877
   w/o Stage1 0.285 0.216 0.095 0.243 0.275 0.244 0.085 0.217 2.841
   w/o Spatial 0.270 0.209 0.029 0.212 0.269 0.237 0.027 0.195 2.701
   w/o ARGI 0.280 0.212 0.029 0.228 0.288 0.238 0.027 0.196 2.796
   w/ Cross-Attn 0.283 0.211 0.029 0.222 0.286 0.240 0.027 0.191 2.778
5. Applications
PianoFlow demonstrates significant potential across various multimodal scenarios within the broader multimedia ecosystem.
AI-Assisted Music Education. PianoFlow enhances pedagogical utility by converting audio into synchronized 3D demonstrations (Xu, 2025; Gan et al., 2024). By providing real-time gestural guidance from performance recordings, our framework offers scalable visual aids for mastering intricate fingerings and bimanual coordination, supplementing traditional instructor-led teaching with high-fidelity visual references.
Interactive Systems and Virtual Livestreaming. Leveraging the AFC mechanism, our framework supports low-latency interactive systems and virtual livestreaming (Yang et al., 2025a; Chu et al., 2025; Yang et al., 2026). It enables digital avatars to respond instantaneously to live musical inputs or audience prompts, ensuring precise synchronization for real-time virtual environments while addressing the latency bottlenecks prevalent in traditional offline-rendered animation.
Controllable Performance Video Synthesis. PianoFlow provides high-fidelity 3D gestural priors to bridge the modality gap in audio-to-video synthesis (Gan et al., 2025; Chen et al., 2025a; Lin et al., 2025; Wang et al., 2025b). By acting as a structural intermediary, these priors anchor downstream generative video models to physical reality, mitigating common artifacts such as unrealistic finger distortions and facilitating the synthesis of photorealistic, temporally coherent digital human performances.
6. Conclusion
We present PianoFlow, a novel two-stage framework for real-time, audio-driven bimanual piano motion generation. Through a cross-modal distillation scheme that leverages MIDI as privileged information during training, our model effectively captures deep harmonic structures while maintaining an efficient audio-only inference pipeline. To facilitate dynamic coordination, we develop the Asymmetric Role-Gated Interaction (ARGI) module to model time-varying bimanual dependencies. Furthermore, our Autoregressive Flow Continuation (AFC) mechanism enables seamless, temporally coherent streaming motion generation for sequences of arbitrary length. Extensive evaluations on the PianoMotion10M dataset demonstrate that PianoFlow establishes a new state-of-the-art in both motion fidelity and structural accuracy, while achieving inference speeds significantly faster than prior methods. Ultimately, PianoFlow offers a highly effective and scalable solution, paving the way for future advancements in digital human animation and real-time interactive virtual performances.